CableVision has been approached by the City of Mirada to run its cable operations in 2019. After negotiating with key parties, CableVision has made the following agreements: It will offer Mirada residents a basic set of 25 cable television stations at a rate of $33.49 per month. . It will pay the city $1,100,000 per year plus $3.00 per cable subscriber per month to maintain the physical facilities. e It will actually pay another company a monthly fixed fee of $55,000 plus $8.50 per cable subscriber per month to broadcast the 25 channels. CableVision estimates that operating costs for billing, program news mailings, etc. will be $130,000 per month plus 8% of monthly revenue. CableVision has several questions about its monthly revenues, costs, and profits in 2019 REQUIRED ROUND YOUR ANSWER TO PART A, QUESTION 1 TO THE NEAREST CENT; ROUND ALL OTHER ANSWERS TO THE NEAREST UNIT OR NEAREST DOLLAR. Part A (8 tries; 8 points) 1. What is the estimated monthy contribution margin per cable subscriber for CableVision in 2019? 300 # 1 IS INCORRECT 2. What are the estimated total monthly fixed costs for CableVision in 2019? #2 IS INCORRECT 55000 Suomit Answer Incorrect. Tries 1/8 Previous Tries Part B (8 tries 8 points) 1. What is CableVisions estimated monthly operating income in 2019 if 18,000 residents subscribe? $70,913 2. How many monthly subscribers would be required for CableVision to break even in 2019? 14,328 3. How many monthly subscribers would be required for CableVision to earn $25,000 per month in 2019? 15,622 4. Assuming a tax rate of 30%, how many monthly subscribers would be required for CableVision to earn $25,000 per month in 2019? 16,177 You are correct. r receipt no. is 152-593Previous Tries Part C (4 tries; 4 points) Some of CableVisions managers are uncertain about their estimate of monthly fixed operating costs. Assuming that 22,000 residents subscribe, how large can monthly fixed operating costs be for CableVision to still earn $25,000 per month in 2019 (ignore taxes)? Incorrect on #7
Answer #1

CableVision's estimated monthly operating income in 2019 if 18,000 residents subscribe = sales - variable cost - Fixed cost

= (31.49*18000) - (3.75 + 8.75 + 2.5192)*18000 - (1300000 + 50000 + 125000)

= - $1178526

2. monthly subscribers would be required for CableVision to break even in 2019 = fixed cost / contribution margin

= (1300000 + 50000 + 125000) / [31.49 - (3.75+8.75+2.5192)]

= 1475000 / 16.4708

= 89552.42

3. monthly subscribers would be required for CableVision to earn $25,000 per month in 2019 =

= (Profit + Fixed cost ) / contribution margin

= ($ 25000 + $ 1475000) / 16.4708

= 91070.26

4) Assuming a tax rate of 35%, what must revenue be in order for CableVision to earn $25,000 per month in 2019 =

= (Profit - Tax + Fixed cost ) / contribution margin

= ($ 25000 - $8750 + $ 1475000) / 16.4708

= 90539.01